OpenStudy (riley):

A liter of milk weighs 1, 032g. The butterfat which it contains to the extent of 4% by volume has a density of 0.865g/cm^3. What is the density of the fat-free skimmed milk?

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (anonymous):

1cm^3=1mL 1L*.04=40mL 40mL*.865g/mL=34.6g 1032g-34.6g=997.4g 997.4g/960mL=1.039g/mL

OpenStudy (riley):

wow thanks O_O